# Heralded Emission Detection in Quantum Dot Solids under Twin-Photon Excitation

### v1

@@ 2 Result
Figure 3: Temperature-dependent HED with SNSPD and APD for sample emission detection (a) Normalized cross-correlation $g^{(2)}(\tau)$ at various temperatures tagging the sample emission with an SNSPD. The bunching peak reflects time correlations governed by the photoluminescence lifetime of the InAs/ZnSe quantum dots (QDs). (b) Normalized $g^{(2)}(\tau)$ of a second QD sample measured with an avalanche photodiode (APD). (c) Data from (b) on a logarithmic scale. (d) Extracted lifetimes as a function of temperature, along with bright–dark model fits. The inset illustrates the model, where G, B, and D denote the ground, bright, and dark state, respectively. (e) Brightness and bunching peak area extracted from normalized $g^{(2)}(\tau)$ data as a function of temperature. The bunching area is fitted with a Bose–Einstein model. (f) External quantum efficiency (EQE) of the twin photon downconversion by the sample at different temperatures, following the trend in brightness.
To demonstrate the HED analysis with these count rates, we used it to quantify temperature-dependent lifetimes. Similar to sample 1, all decay traces consistently display at least two components, independent of temperature. With increasing temperature, the fastest decay component, prominent at low temperatures, gradually diminishes. This trend is not caused by the gold-coated substrate, as control measurements on nanocrystals deposited on plain microscope slides also exhibited the same fastest decay component, as demonstrated for sample 1 (Fig. S5). The initial fast decay component in all $g^{(2)}$ traces is consistent with the rapid decays previously reported in InSb [47, 48] and InAs/CdSe [49] nanocrystals. Such behavior is commonly attributed either to nonradiative recombination [47, 48] or to relaxation from the bright to the dark exciton state [49]. The slower decay components are interpreted in terms of thermally activated dark-to-bright state population [50, 51], corresponding to the average lifetimes of the bright and dark exciton states. The whole lifetimes were extracted using a biexponential model for $T\leq 52.0$ K, where one component captures the initial fast decay and the other represents the average lifetimes of the bright and dark states. At higher temperatures, where the fast component is absent, a single-exponential fit was applied to maintain consistency. This average lifetime can be interpreted as arising from two excitonic states in thermal equilibrium, separated by an energy splitting $\Delta E$. The dark state exhibits a longer lifetime, $\tau_{D}$, whereas the bright state has a shorter lifetime, $\tau_{B}$, in agreement with the bright–dark exciton model.
$\tau^{-1}_{avg}=\left(\frac{e^{\Delta E/k_{B}T}}{1+e^{\Delta E/k_{B}T}}\right)\tau^{-1}_{D}+\left(\frac{1}{1+e^{\Delta E/k_{B}T}}\right)\tau^{-1}_{B}$ (1)
The fitting results (Fig. 3d) yield $\tau_{D}\approx 228.3$ ns, $\tau_{B}\approx 16.6$ ns, and $\Delta E\approx 6.7$ meV, values comparable to those reported for InP/ZnSe QDs [52]. The discrepancies between the data and the fit at 6.3 K, likely arise from either insufficient separation of the fastest decay component during lifetime extraction or from the need for a more comprehensive model that explicitly includes the fast component along with relaxation between bright and dark states or other contributing processes.
Fig. 3e further display the relationship between the sample brightness and the total area under the $g^{(2)}(\tau)$ bunching peak in Fig. 3b, which both diminish with temperature. We point out that the decrease in brightness qualitatively matches a Bose-Einstein model for phonon population with a characteristic energy of $\approx 65~meV$, consistent with phonon energies reported for III–V QDs [52, 53], possibly suggesting phonon-mediated turn-on of non-radiative decay. This effect was less pronounced in Sample 1 (Fig. S9).
@@ 3 Discussion
Advancing quantum spectroscopy will require both improved apparatus capabilities and broader material platforms. This study contributes by identifying colloidal InAs/ZnSe QDs as bright NIR model systems and by detailing heralding analysis, detector tradeoffs, and unexpected artifact sources to guide future improvements. The main limitation in quantum spectroscopy generally is the low coincidence rate after sample interaction [19, 21]. Pinpointing losses to external quantum efficiency across setups [46]and materials is a logical next step. Given the tunability and integrability of QDs as photonic building blocks [54], their targeted optimization could deliver disproportionate gains.
@@
Table 1 compares key parameters and achieved coincidence rates of all HED studies known to the authors, all but Eshun et al. reported low - at most around 120 -coincidence counts per second. One limitation of organic molecules is their generally decreasing luminescent quantum yield (QY) with decreasing emission energy. Few dyes in the near- and shortwave infrared have appreciable QY [39], naturally limiting the coincidence rate for molecules excitable with established sources typically operating in the NIR. Eshun et al. is a notable exception, producing 488–550 nm photons in the sample arm of non-degenerate SPDC pairs, enabling excitation of Rhodamine 6G with an appreciable quantum yield (95%). While this emission efficiency is certainly contributing to the suggested coincidence rate, we point out that there is currently no agreed on best practice for calculating these rates.
Our NIR QDs, used in conjunction with our setup, represent the brightest NIR samples applied to HED to date. We were unable to further improve brightness through systematic thin-film preparation, including multilayer spin-coating and the use of polymeric additives. Importantly, simply maximizing the thin-film PLQY is insufficient to predict the highest HED response, as brightness is also limited by film thickness and morphology. Thicker films can lead to reabsorption losses, while inhomogeneous films may scatter or trap emission, reducing out-coupling efficiency [55]. Thus, optimization of both absorption and emission into Gaussian-like modes, along with minimization of reabsorption in thicker films, is essential to achieve maximum HED brightness.
It is instructive to identify the limiting factors in the downconversion efficiency. Beyond the intrinsic loss from the objective’s collection efficiency, an important factor is the sample configuration—specifically, the number of QDs within the focal volume. Unlike organic dyes, QDs are physically larger, which lowers the probability of absorbing all incident photons, even though they typically exhibit higher PLQY. In addition, when using SNSPDs, the collection efficiency is constrained by the single-mode fibers [46]. While these fibers provide localized information from the film, they also introduce substantial count losses because the emission beam is imaged onto the fiber core, whereas the QD emission across the thin film is spatially inhomogeneous.
@@
In our configuration—using an SNSPD in the heralding arm and a large-area NIR APD in the emission arm—we achieved high coincidence count rates of $\sim$230 coincidence counts per second (ccps) at the low temperatures. This setup maximizes detection efficiency for both herald and emitted photons while avoiding fiber-coupling insertion and losses from the incoherent sample emission into non-Gaussian modes. The trade-off, however, is the APD’s larger timing jitter, which broadens the instrument response function and reduces the precision of $g^{(2)}(\tau)$. These results highlight the importance of detector choice in entangled-photon spectroscopy and suggest that advances in quantum light sources and detectors will further benefit studies of quantum light–matter interactions [15].
@@
Overall, this study establishes colloidal QDs as a solid-state platform for exploring entangled photon light–matter interactions and for developing twin-photon detectors and entangled-photon downconversion devices. While our data do not resolve any effects of prolonged sample coherences on entangled-photon absorption, future phase-sensitive studies—such as Franson-type interferometry [56]—could directly probe entanglement preservation and benefit from the reported brightness herein.
@@
Harper et al. [19] ICG 0.62 CW $1\times 10^{5}$ 860 6.5
Li et al. [21] LH2 1.20 Pulsed $2.7\times 10^{5}$ $9.13\times 10^{5}$ 121
Eshun et al. [20]* R6G 3.99 CW $1.6\times 10^{6}$ 23000
Gäbler et al. [23] IR-140 0.93 CW
Alvarez-Mendoza et al. [22] LH2 1.14 CW $2\times 10^{5}$
@@ 4 Conclusions
We have demonstrated heralded emission detection (HED) under entangled-photon excitation in InAs/ZnSe quantum dot films at cryogenic temperatures, establishing a solid-state platform for this spectroscopy. By benchmarking against TCSPC, we show that HED reliably recovers excitonic lifetimes and bright–dark state dynamics, validating its extension beyond molecular systems. Our results highlight both opportunities and limits. QDs provide brightness, tunability, and device compatibility that make them promising for quantum-light spectroscopy. While this work produced higher coincidence rates than all but one previous work [20], and the highest in the NIR spectral range, the observed coincidence rates are still modest. The demonstrated integration of QDs with optical cavities and waveguides presents a new control knob in quantum light spectroscopy, positioning QDs as a versatile model system.
@@ 5.1 Synthesis of InAs/ZnSe QDs
InAs/ZnSe QDs were synthesized following previously reported methods [57, 58, 59]. Both InAs nanoclusters and InAs QDs were prepared using the same precursor.
@@ 5.2 InAs/ZnSe QDs characterization
The measurements of UV-vis absorption were performed on a Jasco V-570 UV-vis-NIR spectrophotometer. Photoluminescence spectra and photoluminescence quantum yield was measured by a Hamamatsu Absolute PL quantum yield spectrometer C11347-11 and Edinburgh FLS920 fluorescence spectrometer. Transmission electron microscopy (TEM) images of the InAs/ZnSe QDs were acquired using a Tecnai 12 TEM. Scanning electron microscopy (SEM) images of the QDs film were obtained with a Zeiss XB 550 FIB-SEM.
@@ Acknowledgments
We thank Tim Rambo and Ryan Wilson for helpful discussions on instrumentation. This work was financially supported by the College of Chemistry at the University of California, Berkeley and the US Department of Energy (DE-AC02-05CH11231).
@@
The entangled photon pair source (customized QES 2.4, Qubitekk) consists of a 405 nm continuous-wave (CW) laser that pumps a periodically poled potassium titanyl phosphate (ppKTP) crystal housed in a temperature-controlled oven, allowing for tuning of the photon wavelength and degree of degeneracy. The indistinguishability of photon pairs has been shown in Fig. S1. Entangled photon pairs are generated via spontaneous parametric down-conversion (SPDC), after which the residual pump light is filtered out. The entangled photons are then coupled into optical fiber. A polarization-maintaining, fiber-coupled polarizing beam splitter (PBS) with FC/PC connectors (Qubitekk) is connected to the source output to separate the signal and idler photons. One output of the PBS is routed via an FC/PC sleeve (ADAFC2, Thorlabs) to a single-mode fiber patch cable (780HP, Diamond), which is terminated with an E2000 connector and connected to a superconducting nanowire single-photon detector (SNSPD, Quantum Opus), referred to as the heralding arm.
@@
Detection events from both SNSPDs are time-stamped with 8 ps resolution using a single-photon counting module (Time Tagger Ultra, Performance Edition, Swabian Instruments). For second-order correlation measurements, the system accumulates the time differences between photon detection events on the two channels into a histogram. Each detection event is considered as both a “start” and a “stop,” allowing the calculation of both positive and negative time delays. A time bin width of 1 ns is chosen to balance the trade-off between signal-to-noise ratio (SNR) and timing resolution.
For the APD-based version of the HED, the short-pass dichroic mirror (FF825-SDI01, Semrock) replaces the PBS output path, and the signal is detected using an APD (id120, IDQ).
@@
Figure S1: Hong–Ou–Mandel (HOM) dip of the entangled photon pairs, indicating their indistinguishability, with a visibility of approximately 86%.
@@
The setup for TCSPC measurements closely follows that used for HED, with the primary difference being the excitation source: a 470 nm pulsed laser (LDH-P-C-470, PicoQuant). The laser repetition rate is reduced to 2.5 MHz using an external controller (PDL 820, PicoQuant) to meet the desired excitation conditions. Along the excitation path, a 500 nm short-pass filter (FESH0500, Thorlabs) is combined with a set of neutral density filters (NE20B, NE20A, and NE10A-A; Thorlabs), reducing the excitation fluence to 6.37 nJ/cm², which cause the exciton number per pulse $\langle N\rangle\ll 1$. All other optical components remain unchanged from the HES setup.

### v2

@@ 1 Result
Figure 3: Temperature-dependent HED with SNSPD and APD for sample emission detection. (a) Normalized cross-correlation, $g^{(2)}(\tau)$, at various temperatures, measured by tagging the sample emission with an SNSPD. The bunching peak reflects the time correlations determined by the photoluminescence lifetime of the InAs/ZnSe quantum dots (QDs). The inset shows a zoomed-in view on a logarithmic scale, with $g^{(2)}(\tau)$ normalized so that all peak values are equal, facilitating comparison of the decay dynamics. (b) $g^{(2)}(\tau)$ of a second QD sample measured with an avalanche photodiode (APD). (c) Data from (b) plotted as $G^{(2)}(\tau)$ normalized on a logarithmic scale. (d) Extracted lifetimes as a function of temperature, along with bright–dark model fits. The inset illustrates the model, where G, B, and D denote the ground, bright, and dark state, respectively. (e) Brightness and bunching peak area extracted from $g^{(2)}(\tau)$ data as a function of temperature. The bunching area is fitted with a Bose–Einstein model. (f) External quantum efficiency (EQE) of the twin photon downconversion by the sample at different temperatures, following the trend in brightness.
To demonstrate the HED analysis under these count rates, we applied it to quantify temperature-dependent lifetimes. The decay dynamics at all temperatures are well described by exponential models containing at least two components, consistent with Sample 1. Specifically, for temperatures $T\leq 52.0$ K, the data were fitted using a biexponential model, where one component captures an initial fast decay and the other represents the slower average lifetime associated with the bright and dark exciton states. At higher temperatures, where the fast component vanishes, a single-exponential fit was sufficient for consistency.
With increasing temperature, the fastest decay component—dominant at low temperatures—gradually diminishes. This behavior is intrinsic to the nanocrystals and not influenced by the gold-coated substrate, as control samples deposited on plain glass slides exhibit the same feature (see Fig. S9). The presence of this initial fast decay in all HED traces is consistent with previously reported rapid components in time-resolved photoluminescence measurements of InSb [liu2012colloidal, busatto2020luminescent] and InAs/CdSe [bischof2015origin] nanocrystals. Such fast decays are typically attributed to either nonradiative recombination [liu2012colloidal, busatto2020luminescent] or relaxation from the bright to dark exciton state [bischof2015origin]. Analysis of the temperature dependence of the fast component’s fractional amplitude (see Supporting Information) reveals two activation energies, 2.96 meV and 15.99 meV, which can be ascribed respectively to acoustic-phonon–mediated relaxation or shallow surface/ligand traps, and to deeper trap states or defect-assisted nonradiative decay.
The slower decay component is attributed to thermally activated dark-to-bright exciton population [de2006size, oron2009universal], reflecting the effective lifetime of excitons in thermal equilibrium. The average lifetime extracted from these fits can be modeled as arising from two thermally coupled excitonic states separated by an energy splitting $\Delta E$, together with an additional trap-assisted decay channel [murphy2016temperature, gaponenko2010temperature]. The bright state has a shorter lifetime $\tau_{B}$, while the dark state has a longer lifetime $\tau_{D}$, consistent with the bright–dark exciton framework [murphy2016temperature, gaponenko2010temperature]:
$\tau_{\mathrm{avg}}^{-1}=\frac{a\tau_{B}^{-1}+\tau_{D}^{-1}+k_{0}\left[a(b-1)^{-n}+(b-1)^{-m}\right]}{1+a},$ (1)
where
$a=e^{-\Delta E/(k_{B}T)},\quad b=e^{E_{\mathrm{ph}}/(k_{B}T)}.$ (2)
Here, $k_{0}$ denotes the trap-assisted decay rate, $E_{\mathrm{ph}}$ the relevant phonon energy, and $n$ and $m$ the phonon interaction exponents ($m>n$).
Fitting results (Fig. 3d) yield $\tau_{D}\approx 246.6$ ns, $\tau_{B}\approx 21.4$ ns, $\Delta E\approx 5.5$ meV, $E_{\mathrm{ph}}\approx 23.8$ meV, $1/k_{0}\approx 1.03$ ns, $m\approx 1.03$, and $n\approx 0.79$. The extracted $\tau_{D}$ is shorter than the microsecond-scale dark-state lifetimes typically observed in InAs QDs [oron2009universal], but comparable to those of InP/ZnSe QDs [brodu2018exciton]. The obtained $\Delta E$ agrees well with reported bright–dark energy splittings of 2–9 meV in other III–V systems [brodu2018exciton, chandrasekaran2023exciton]. The derived phonon energy $E_{\mathrm{ph}}$ also aligns with optical phonon energies (24–32 meV) of the ZnSe shell [brodu2018exciton]. The small deviations observed at 6.3 K may result from incomplete isolation of the fastest decay component during fitting or the need for a more comprehensive model that explicitly includes fast relaxation dynamics between bright and dark states.
Finally, Fig. 3e illustrates how both the sample brightness and the total area under the $g^{(2)}(\tau)$ bunching peak (Fig. 3b) decrease with increasing temperature. The reduction in brightness follows a Bose–Einstein–like dependence on phonon population, with a fitted characteristic energy of $\approx 65$ meV—consistent with phonon-mediated activation of nonradiative decay in III–V QDs [brodu2018exciton, brodu2020exciton]. The reduced bunching peak area can be explained by the increased influence of detector dark counts as QD emission weakens at higher temperature, leading to a lower signal-to-dark-count ratio and diminished bunching contrast. This effect is less pronounced in Sample 1, where SNSPDs with lower dark count rates were used.
@@ 2 Discussion
Advancing quantum spectroscopy will require both improved apparatus capabilities and broader material platforms. This study contributes by identifying colloidal InAs/ZnSe QDs as bright NIR model systems and by detailing heralding analysis, detector tradeoffs, and unexpected artifact sources to guide future improvements. The main limitation in quantum spectroscopy generally is the low coincidence rate after sample interaction [harper2023entangled, li2023single]. Pinpointing losses to external quantum efficiency across setups [kim2011efficient] and materials is a logical next step. Given the tunability and integrability of QDs as photonic building blocks [ibanez2025prospects], their targeted optimization could deliver disproportionate gains.
@@
Table 1 compares key parameters and achieved coincidence rates of all HED studies known to the authors, all but Eshun et al. reported low - at most around 120 -coincidence counts per second. One limitation of organic molecules is their generally decreasing luminescent quantum yield (QY) with decreasing emission energy. Few dyes in the near- and shortwave infrared have appreciable QY [rurack2011fluorescence], naturally limiting the coincidence rate for molecules excitable with established sources typically operating in the NIR. Eshun et al. is a notable exception, producing 488–550 nm photons in the sample arm of non-degenerate SPDC pairs, enabling excitation of Rhodamine 6G with an appreciable quantum yield (95%). While this emission efficiency is certainly contributing to the suggested coincidence rate, we point out that there is currently no agreed-upon best practice for calculating these rates.
Our NIR QDs, when integrated with our detection setup, represent the brightest NIR emitters applied to HED to date. However, maximizing the thin-film PLQY alone is insufficient to achieve the highest HED response, as the detected brightness is also influenced by film thickness, morphology, and sample configuration. Thicker films can lead to reabsorption losses, while inhomogeneous films may cause scattering or trapping of emission, reducing photon out-coupling. In addition, the number of QDs within the focal volume plays a critical role: unlike smaller organic dyes, the larger physical size of QDs lowers the probability of absorbing all incident photons, even though their PLQY is typically higher. Detection using single-mode fiber–coupled SNSPDs introduces further constraints. While these fibers provide spatially localized information from the film, they inherently limit collection efficiency because the QD emission is spatially inhomogeneous across the thin film [li2018effect, kim2011efficient]. Therefore, achieving optimal HED brightness requires simultaneous optimization of absorption and emission into Gaussian-like optical modes, minimizing reabsorption in thicker films, and carefully considering QD density and emission distribution. These considerations are crucial for enhancing photon out-coupling and extending the applicability of HED in quantum photonic systems.
@@
In our other configuration—using an SNSPD in the heralding arm and a large-area NIR APD in the emission arm—we achieved high coincidence count rates of $\sim$230 coincidence counts per second (ccps) at low temperatures. This setup maximizes detection efficiency for both herald and emitted photons while avoiding fiber-coupling insertion and losses from the incoherent sample emission into non-Gaussian modes. The trade-off, however, is the APD’s larger timing jitter, which broadens the instrument response function (Fig. S1b) and reduces the precision of $g^{(2)}(\tau)$. These results highlight the importance of detector choice in entangled-photon spectroscopy and suggest that advances in quantum light sources and detectors will further benefit studies of quantum light–matter interactions [tsao2025enhancing].
@@
Overall, this study establishes colloidal QDs as a versatile solid-state platform for exploring entangled-photon light–matter interactions and for developing twin-photon detectors and entangled-photon downconversion devices. Furthermore, it provides a reliable approach to study photon-bunching dynamics modulated by emitter lifetimes. While our current data do not resolve effects of prolonged sample coherences on entangled-photon absorption, future phase-sensitive experiments—such as Franson-type interferometry [fasel2005energy]—could directly probe entanglement preservation and coherence dynamics. The high brightness of our NIR QDs demonstrated here not only facilitates such measurements but also opens avenues for their integration into practical quantum photonic systems, including quantum communication, quantum metrology, and on-chip entangled-photon sources. Taken together, these results highlight the promise of colloidal QDs as both a fundamental platform for studying quantum light–matter interactions and a practical building block for emerging quantum technologies.
@@
Harper et al. [harper2023entangled] ICG 0.62 CW $1\times 10^{5}$ 860 6.5
Li et al. [li2023single] LH2 1.20 Pulsed $2.7\times 10^{5}$ $9.13\times 10^{5}$ 121
Eshun et al. [eshun2023fluorescence]* R6G 3.99 CW $1.6\times 10^{6}$ 23000
Gäbler et al. [gabler2025benchmarking] IR-140 0.93 CW
Alvarez-Mendoza et al. [alvarez2025correlated] LH2 1.14 CW $2\times 10^{5}$
@@ 3 Conclusions
We have demonstrated heralded emission detection (HED) under entangled-photon excitation in InAs/ZnSe quantum dot films at cryogenic temperatures, establishing a solid-state platform for this spectroscopy. By benchmarking against TCSPC, we show that HED reliably recovers excitonic lifetimes and bright–dark state dynamics, validating its extension beyond molecular systems. Our results highlight both opportunities and limits. QDs provide brightness, tunability, and device compatibility that make them promising for quantum-light spectroscopy. While this work produced higher coincidence rates than all but one previous work [eshun2023fluorescence], and the highest in the NIR spectral range, the observed coincidence rates are still modest. The demonstrated integration of QDs with optical cavities and waveguides presents a new control knob in quantum light spectroscopy, positioning QDs as a versatile model system.
@@ 4.1 Synthesis of InAs/ZnSe QDs
InAs/ZnSe QDs were synthesized following previously reported methods [asor2023zn, li2025shell, tamang2016tuning]. Both InAs nanoclusters and InAs QDs were prepared using the same precursor.
@@ 4.2 InAs/ZnSe QDs characterization
The measurements of UV-vis absorption were performed on a Jasco V-570 UV-vis-NIR spectrophotometer. Photoluminescence spectra and photoluminescence quantum yield was measured by a Hamamatsu Absolute PL quantum yield spectrometer C11347-11 and Edinburgh FLS920 fluorescence spectrometer. Transmission electron microscopy (TEM) images of the InAs/ZnSe QDs were acquired using a Tecnai 12 TEM. Scanning electron microscopy (SEM) images of the QD films were acquired using a Zeiss XB 550 FIB-SEM at 0.7 kV with a working distance of 4 mm.
@@ Acknowledgments
We thank Tim Rambo and Ryan Wilson for valuable discussions regarding instrumentation. This work was supported by the College of Chemistry at the University of California, Berkeley. Work at the Molecular Foundry was supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. Partial support was also provided by the Israel Science Foundation through the MAPATS program (U.B., Grant No. 2655/23).
@@
The entangled photon pair source (customized QES 2.4, Qubitekk) consists of a 405 nm continuous-wave (CW) laser that pumps a periodically poled potassium titanyl phosphate (ppKTP) crystal housed in a temperature-controlled oven, allowing for tuning of the photon wavelength and degree of degeneracy. The indistinguishability of photon pairs has been shown in Fig. S1a. Entangled photon pairs are generated via spontaneous parametric down-conversion (SPDC), after which the residual pump light is filtered out. The entangled photons are then coupled into an optical fiber. A polarization-maintaining, fiber-coupled polarizing beam splitter (PBS) with FC/PC connectors (Qubitekk) is connected to the source output to separate the signal and idler photons. One output of the PBS is routed via an FC/PC sleeve (ADAFC2, Thorlabs) to a single-mode fiber patch cable (780HP, Diamond), which is terminated with an E2000 connector and connected to a superconducting nanowire single-photon detector (SNSPD, Quantum Opus), referred to as the heralding arm.
@@
Detection events from both SNSPDs are time-stamped with 19 ps (8 ps in RMS) resolution using a single-photon counting module (Time Tagger Ultra, Performance Edition, Swabian Instruments). For second-order correlation measurements, the system accumulates the time differences between photon detection events on the two channels into a histogram. Each detection event is considered as both a “start” and a “stop,” allowing the calculation of both positive and negative time delays. A time bin width of 1 ns is chosen to balance the trade-off between signal-to-noise ratio (SNR) and timing resolution.
For the APD-based version of the HED, the short-pass dichroic mirror (FF825-SDI01, Semrock) replaces the PBS output path, and the signal is detected using an APD (id120, IDQ). The impulse response function (IRF) for this setup is shown in Fig. S1b.
@@
The setup for TCSPC measurements closely follows that used for HED, with the primary difference being the excitation source: a 470 nm pulsed laser (LDH-P-C-470, PicoQuant). The laser repetition rate is reduced to 2.5 MHz using an external controller (PDL 820, PicoQuant) to meet the desired excitation conditions. Along the excitation path, a 500 nm short-pass filter (FESH0500, Thorlabs) is combined with a set of neutral density filters (NE20B, NE20A, and NE10A-A; Thorlabs), reducing the excitation fluence to 6.37 nJ/cm², which cause the exciton number per pulse $\langle N\rangle\ll 1$ (See the following section to discuss the value). All other optical components remain unchanged from the HED setup.